Ferrell Inc. recently reported net income of $9 million. It has 630,000 shares of common stock, which currently trades at $29 a share. Ferrell continues to expand and anticipates that 1 year from now, its net income will be $13.05 million. Over the next year, it also anticipates issuing an additional 220,500 shares of stock so that 1 year from now it will have 850,500 shares of common stock. Assuming Ferrell's price/earnings ratio remains at its current level, what will be its stock price 1 year from now? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Please provide rating... let me know if you need any clarification..

Computation of current PE ratio
Earning $       9,000,000
Number of share 630000
EPS                  14.29
Share price $              29.00
P/E ratio=29/14.29                    2.03
Computation of after 1 year share price (Assuming same PE ratio)
i Earning         13,050,000
ii Number of share 850,500
iii=i/ii Earning per share                  15.34
iv PE ratio                    2.03
v=iii*iv New share price                  31.15
Hence, new share price = $              31.15